The first sample shows that math works. I let the line be drawn to make the sample clearer.

\large
\pstextpath[c]%
{\pscurve[linecolor=gray](0,1)(4,3)(6,2)(9,0)(12,1)(15,1)}%
{$S_\alpha=\Omega(\gamma_\beta)$ is a connected snarf and
$B=(\otimes,\rightarrow,\theta)$ is Boolean left subideal.}

S_{α} = Ω(γ_{β}) is a connected snarf and B = (⊗, →, θ) is Boolean left subideal.

\psset{linestyle=none}
\pstextpath[c]{\psarcn(0,0){73pt}{180}{0}}%
{Centre National de la}
\pstextpath[c]{\psarc(0,0){73pt}{180}{0}}%
{Recherche Scientifique}

Centre National de la Recherche Scientifique

In principle, it is possible to use parbox, but let’s see what really happens. It seems hard to believe that someone would want to do this.